OpenStudy (anonymous):

What is the simplified form of 4z^2-16z+15/2z^2+z-15

OpenStudy (mertsj):

Factor the numerator and denominator and cancel any common factors.

OpenStudy (anonymous):

(z-6)(z-10)/(z+6)(z-5)?

OpenStudy (mertsj):

What is (z-6)(z-10)?

OpenStudy (anonymous):

factor of the numerator?

OpenStudy (anonymous):

oh snap, haha sorry

OpenStudy (anonymous):

oh no, its 4*15=60. -6 - 10= -16

OpenStudy (mertsj):

\[\frac{(2z-5)(2z-3)}{(2z-5)(z+3)}\]